OpenStudy (anonymous):

Write the equations for the lines described 1) containing the points (1,-3) and (5,-2) For this question I actually solved it and attained y = -5/6 + 13/6. I just need someone to verify these for me. 2) with slope 0 and containing the point (4,2)

OpenStudy (shamil98):

(1,-3) and (5,-2) slope = -2 -(-3) / 5 - 1 = 1 /4 use the point-slope formula. y - (-3) = 1/4(x-1) y + 3 = 1/4x - 1/4 y = 1/4x - 13/4

OpenStudy (shamil98):

2) m = 0 (4,2) y - 2 = 0(x-4) y - 2= 0x + 0 y - 2 =0 y = 2